# SPIdepth: Strengthened Pose Information for Self-supervised Monocular Depth Estimation

### v1

@@ Abstract
Self-supervised monocular depth estimation has garnered considerable attention for its applications in autonomous driving and robotics. While recent methods have made strides in leveraging techniques like the Self Query Layer (SQL) to infer depth from motion, they often overlook the potential of strengthening pose information. In this paper, we introduce SPIdepth, a novel approach that prioritizes enhancing the pose network for improved depth estimation. Building upon the foundation laid by SQL, SPIdepth emphasizes the importance of pose information in capturing fine-grained scene structures. By enhancing the pose network’s capabilities, SPIdepth achieves remarkable advancements in scene understanding and depth estimation. Experimental results on benchmark datasets such as KITTI and Cityscapes showcase SPIdepth’s state-of-the-art performance, surpassing previous methods by significant margins. Notably, SPIdepth’s performance exceeds that of unsupervised models and, after finetuning on metric data, outperforms all existing methods. Remarkably, SPIdepth achieves these results using only a single image for inference, surpassing even methods that utilize video sequences for inference, thus demonstrating its efficacy and efficiency in real-world applications. Our approach represents a significant leap forward in self-supervised monocular depth estimation, underscoring the importance of strengthening pose information for advancing scene understanding in real-world applications.
@@ 1 Introduction
SPIdepth extends the capabilities of SQL by strengthened robust pose information, which is crucial for interpreting complex spatial relationships within a scene. Our extensive evaluations on benchmark datasets such as KITTI and Cityscapes demonstrate SPIdepth’s superior performance, surpassing previous self-supervised methods in both accuracy and generalization capabilities. Remarkably, SPIdepth achieves these results using only a single image for inference, outperforming methods that rely on video sequences.
@@
  • Our method sets a new benchmark in self-supervised depth estimation, outperforming existing methods on standard datasets like KITTI and Cityscapes, using only a single image for inference, without the need for video sequences.
@@ 3 Methodology
We address the task of self-supervised monocular depth estimation, focusing on predicting depth maps from single RGB images without ground truth, akin to learning structure from motion (SfM). Our approach, SPIdepth, introduces strengthen the pose network and enhances depth estimation accuracy. Unlike conventional methodologies that primarily focus on depth refinement, SPIdepth prioritizes improving the accuracy of the pose network to capture intricate scene structures more effectively, leading to significant advancements in depth prediction accuracy.
@@
Table 2: Comparison with supervised methods on KITTI [13] eigen benchmark using self-supervised pretrained and metric fine-tuned model. Scaling for using median-scaling to get absolute depth scale.
@@ 4 Results
Our assessment of SMIDepth encompasses two widely-used datasets: KITTI and Cityscapes, employing established evaluation metrics.
@@ 4.2 KITTI Results
We present the performance comparison of SMIDepth with several state-of-the-art self-supervised depth estimation models on the KITTI dataset, as summarized in Table 1. SMIDepth achieves superior performance compared to all other models across various evaluation metrics. Notably, it achieves the lowest values of AbsRel (absolute relative difference), SqRel (squared relative difference), RMSE (root mean squared error), and RMSElog (logarithmic RMSE), indicating its exceptional accuracy in predicting depth values.
In particular, SMIDepth demonstrates remarkable accuracy even when compared to models like ManyDepth, which leverage multiple frames for testing. This indicates SMIDepth’s ability to achieve state-of-the-art results using only a single frame as input during testing, highlighting its efficiency and practicality.
@@
Moving on to Table 2, we compare the performance of SMIDepth with several supervised depth estimation models on the KITTI eigen benchmark. Despite being self-supervised and metric fine-tuned, SMIDepth outperforms supervised methods in terms of AbsRel, SqRel, RMSE, and RMSElog. Notably, SMIDepth achieves the lowest values across all these metrics, indicating its superior accuracy in predicting depth values.
@@
Furthermore, SMIDepth surpasses LightedDepth, a model that operates on video sequences (more than one frame) and outperforms a good pre-trained models like EVP based on stable diffusion [30]. Despite LightedDepth’s advantage of using multiple frames, SMIDepth still outperforms it, highlighting SMIDepth’s robustness and effectiveness even in challenging scenarios.
@@
Additionally, SMIDepth demonstrates significant performance improvements over SQLdepth, a model that serves as the foundation for its development. Notably, SMIDepth outperforms SQLdepth by a large margin in both supervised and self-supervised settings. This substantial improvement underscores the impact of strengthening the pose net and its information in SMIDepth.
@@
Overall, these results underscore the effectiveness of SMIDepth in self-supervised monocular depth estimation, positioning it as a leading model in the field.
@@ 5 Conclusion
In summary, our method, SPIdepth, stands out for its significant advancements in self-supervised monocular depth estimation. By enhancing the pose network and leveraging ConvNext for depth prediction, we achieve exceptional accuracy. Across KITTI and Cityscapes datasets, SPIdepth demonstrates versatility and effectiveness, even in scenarios with dynamic objects. Notably, on KITTI, our approach sets new benchmarks in both self-supervised training and subsequent fine-tuning, surpassing models using multiple frames per inference. These results underscore SPIdepth’s potential for real-world applications requiring precise depth estimation.

### v2

@@ Abstract
Self-supervised monocular depth estimation has garnered considerable attention for its applications in autonomous driving and robotics. While recent methods have made strides in leveraging techniques like the Self Query Layer (SQL) to infer depth from motion, they often overlook the potential of strengthening pose information. In this paper, we introduce SPIdepth, a novel approach that prioritizes enhancing the pose network for improved depth estimation. Building upon the foundation laid by SQL, SPIdepth emphasizes the importance of pose information in capturing fine-grained scene structures. By enhancing the pose network’s capabilities, SPIdepth achieves remarkable advancements in scene understanding and depth estimation. Experimental results on benchmark datasets such as KITTI, Cityscapes, and Make3D showcase SPIdepth’s state-of-the-art performance, surpassing previous methods by significant margins. Specifically, SPIdepth tops the self-supervised KITTI benchmark. Additionally, SPIdepth achieves the lowest AbsRel (0.029), SqRel (0.069), and RMSE (1.394) on KITTI, establishing new state-of-the-art results. On Cityscapes, SPIdepth shows improvements over SQLdepth of 21.7% in AbsRel, 36.8% in SqRel, and 16.5% in RMSE, even without using motion masks. On Make3D, SPIdepth in zero-shot outperforms all other models. Remarkably, SPIdepth achieves these results using only a single image for inference, surpassing even methods that utilize video sequences for inference, thus demonstrating its efficacy and efficiency in real-world applications. Our approach represents a significant leap forward in self-supervised monocular depth estimation, underscoring the importance of strengthening pose information for advancing scene understanding in real-world applications.
@@ 1 Introduction
SPIdepth extends the capabilities of SQL by strengthened robust pose information, which is crucial for interpreting complex spatial relationships within a scene. Our extensive evaluations on benchmark datasets such as KITTI, Cityscapes, Make3D and demonstrate SPIdepth’s superior performance, surpassing previous self-supervised methods in both accuracy and generalization capabilities. Remarkably, SPIdepth achieves these results using only a single image for inference, outperforming methods that rely on video sequences. Specifically, SPIdepth tops the self-supervised KITTI benchmark. Additionally, SPIdepth achieves the lowest AbsRel (0.029), SqRel (0.069), and RMSE (1.394) on KITTI, establishing new state-of-the-art results. On Cityscapes, SPIdepth shows improvements over SQLdepth of 21.7% in AbsRel, 36.8% in SqRel, and 16.5% in RMSE, even without using motion masks. On Make3D, SPIdepth in zero-shot outperforms all other models.
@@
  • Our self-supervised method sets a new benchmark in depth estimation, outperforming all existing methods on standard datasets like KITTI and Cityscapes using only a single image for inference, without the need for video sequences. Additionally, our approach achieves significant improvements in zero-shot performance on the Make3D dataset.
@@ 2.3 Self-supervised Depth Estimation
Figure 1: The SPIdepth architecture. An encoder-decoder extracts features from frame $I_{t}$, which are then input into the Self Query Layer to obtain the depth map $D_{t}$. Strengthened PoseNet predicts the relative pose between frame $I_{t}$ and reference frame $I_{t}^{\prime}$ using a powerful pose network, needed only during training. Pixels from frame $I_{t}^{\prime}$ are used to reconstruct frame $I_{t}$ with depth map $D_{t}$ and relative pose $T_{t^{\prime}\rightarrow t}$. The loss function is based on the differences between the warped image $I_{t^{\prime}\rightarrow t}$ and the source image $I_{t}$.
@@ 3 Methodology
We address the task of self-supervised monocular depth estimation, focusing on predicting depth maps from single RGB images without ground truth, akin to learning structure from motion (SfM). Our approach, SPIdepth  fig. 1, introduces strengthen the pose network and enhances depth estimation accuracy. Unlike conventional methodologies that primarily focus on depth refinement, SPIdepth prioritizes improving the accuracy of the pose network to capture intricate scene structures more effectively, leading to significant advancements in depth prediction accuracy.
@@
Table 2: Comparison with supervised methods on KITTI [14] eigen benchmark using self-supervised pretrained and metric fine-tuned model. The best results are in bold, and second best are underlined.
@@ 4 Results
Our assessment of SMIDepth encompasses three widely-used datasets: KITT, Cityscapes and Make3D, employing established evaluation metrics.
@@ 4.1.2 Cityscapes Dataset
4.1.3 Make3D Dataset
Make3D [38] is a monocular depth estimation dataset containing 400 high-resolution RGB and low-resolution depth map pairs for training, and 134 test samples. To evaluate SMIDepth’s generalization ability on unseen data, zero-shot evaluation on the Make3D test set has been performed using the SMIDepth model pre-trained on KITTI.
@@ 4.2 KITTI Results
We present the performance comparison of SMIDepth with several state-of-the-art self-supervised depth estimation models on the KITTI dataset, as summarized in Table 1. SMIDepth achieves superior performance compared to all other models across various evaluation metrics. Notably, it achieves the lowest values of AbsRel (0.071), SqRel (0.531), RMSE (3.662), and RMSElog (0.153), indicating its exceptional accuracy in predicting depth values.
@@
Moving on to Table 2, we compare the performance of SMIDepth with several supervised depth estimation models on the KITTI eigen benchmark. Despite being self-supervised and metric fine-tuned, SMIDepth outperforms supervised methods across all these metrics, indicating its superior accuracy in predicting metric depth values.
@@
Furthermore, SMIDepth surpasses LightedDepth, a model that operates on video sequences (more than one frame) and outperforms a good pre-trained models like EVP based on stable diffusion [36]. Despite LightedDepth’s advantage of using multiple frames, SMIDepth shows improvements of 0.012 (29.3%) in AbsRel, 0.038 (34.3%) in SqRel, 0.354 (20.3%) in RMSE, and 0.011 (18.6%) in RMSElog, highlighting SMIDepth’s robustness and effectiveness even in challenging scenarios.
@@
Additionally, SMIDepth demonstrates significant performance improvements over SQLdepth, a model that serves as the foundation for its development. In the self-supervised setting, SMIDepth shows improvements of 5.3% in AbsRel, 1.5% in SqRel, 1.6% in RMSE, and 1.9% in RMSElog. In the supervised setting, SMIDepth shows improvements of 32.6% in AbsRel, 35.6% in SqRel, 17.9% in RMSE, and 25% in RMSElog. These substantial improvements underscore the impact of strengthening the pose net and its information in SMIDepth.
@@
Overall, these results underscore the effectiveness of SMIDepth in self-supervised monocular depth estimation, positioning it as a leading model in the field. Qualitative results further illustrate the superior performance of SMIDepth, as shown in Figure 2.
Figure 2: Qualitative results on the KITTI dataset. From left to right: Input RGB image, Ground Truth, SQLdepth prediction, and SPIdepth prediction.
Method Train $AbsRel\downarrow$ $SqRel\downarrow$ $RMSE\downarrow$ $RMESlog\downarrow$ $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$
Pilzer et al. [33] GAN, C 0.240 4.264 8.049 0.334 0.710 0.871 0.937
Struct2Depth 2 [5] MMask, C 0.145 1.737 7.280 0.205 0.813 0.942 0.976
Monodepth2 [16] –, C 0.129 1.569 6.876 0.187 0.849 0.957 0.983
Videos in the Wild [17] MMask, C 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li et al. [27] MMask, C 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Lee et al. [26] MMask, C 0.116 1.213 6.695 0.186 0.852 0.951 0.982
ManyDepth [48] MMask, C 0.114 1.193 6.223 0.170 0.875 0.967 0.989
InstaDM [25] MMask, C 0.111 1.158 6.437 0.182 0.868 0.961 0.983
[…]
ProDepth [49] MMask, C 0.095 0.876 5.531 0.146 0.908 0.978 0.993
RM-Depth [19] MMask, C 0.090 0.825 5.503 0.143 0.913 0.980 0.993
[…]
4.3 Cityscapes Results
To evaluate the generalization of SMIDepth, we conducted fine-tuning experiments in a self-supervised manner without using a motion mask on the Cityscapes dataset. Starting from a KITTI pre-trained model, we fine-tuned it on Cityscapes. The results, summarized in Table 3, demonstrate that SMIDepth outperforms all other methods, including those that use motion masks.
Despite not using a motion mask—a technique commonly employed to handle the high proportion of moving objects in the Cityscapes dataset—SMIDepth achieves remarkable improvements over other models. Compared to SQLdepth, SMIDepth shows significant advancements: improvements of 0.023 (21.7%) in AbsRel, 0.432 (36.8%) in SqRel, and 1.032 (16.5%) in RMSE.
Moreover, compared to the previous state-of-the-art model RM-Depth, which also uses motion masks, SMIDepth achieves improvements of 0.007 (7.8%) in AbsRel, 0.084 (10.2%) in SqRel, and 0.298 (5.4%) in RMSE.
These results underscore SMIDepth’s exceptional generalization and accuracy, achieved without the use of motion masks. This makes SMIDepth a highly robust and efficient option for depth estimation tasks. Its performance demonstrates its capability for quick deployment in new datasets, effectively addressing the challenges posed by moving objects.
4.4 Make3D results
To assess the generalization capacity of SMIDepth, a zero-shot evaluation was performed on the Make3D dataset [38] using pretrained weights from KITTI. Adhering to the evaluation settings of [15, 45], SMIDepth achieved superior results compared to other methods, including SQLdepth. Table 4 highlights these findings, showcasing the remarkable zero-shot generalization ability of the SMIDepth model.
As summarized in Table 4, SMIDepth achieves the lowest values in all evaluation metrics, with AbsRel (0.299), SqRel (1.931), RMSE (6.672), and $log_{10}$ (0.144). These results highlight the remarkable zero-shot generalization ability of the SMIDepth model, significantly outperforming the previous best model, SQLdepth. The improvements of SMIDepth over SQLdepth are 0.007 (2.3%) in AbsRel, 0.471 (19.6%) in SqRel, 0.184 (2.7%) in RMSE, and 0.007 (4.6%) in $log_{10}$, underscoring its superior performance in challenging zero-shot scenarios.
Method Type $AbsRel\downarrow$ $SqRel\downarrow$ $RMSE\downarrow$ $log_{10}\downarrow$
Monodepth [15] S 0.544 10.94 11.760 0.193
Zhou [62] M 0.383 5.321 10.470 0.478
DDVO [43] M 0.387 4.720 8.090 0.204
Monodepth2 [16] M 0.322 3.589 7.417 0.163
CADepthNet [51] M 0.312 3.086 7.066 0.159
SQLdepth [45] M 0.306 2.402 6.856 0.151
SMIDepth M 0.299 1.931 6.672 0.144
Table 4: Performance comparison on Make3D dataset [38]. The best results are in bold, and second best are underlined.
5 Ablation Study
To assess the impact of Strengthened Pose Information (SPI) on depth estimation performance, we conducted an ablation study using various backbone networks, evaluated on the KITTI dataset in both self-supervised and supervised fine-tuning settings. This study compared ConvNeXt Large, ConvNeXt X-Large, and ConvNeXtV2 Huge with and without SPI, as summarized in Table 5.
Initially, we evaluated ConvNeXt Large using the standard pose net configuration, as employed in previous state-of-the-art approach - SQLdepth [45]. Without SPI, it achieved an AbsRel of 0.075 and RMSE of 3.722 in self-supervised settings. Introducing SPI improved these metrics to AbsRel 0.072 and RMSE 3.677. In supervised fine-tuning, the SPI-enhanced model showed a reduction of 0.006 (14%) in AbsRel, and a reduction of 0.101 (5.9%) in RMSE. ConvNeXt X-Large and ConvNeXtV2 Huge with SPI further improved performance, reaching AbsRel 0.071 and RMSE 3.662 in self-supervised settings, and AbsRel 0.029 and RMSE 1.394 in supervised fine-tuning.
While changing the backbone size provides only slight improvements in the self-supervised setting compared to the impact of SPI, it does result in more significant gains in supervised settings. These results highlight that SPI significantly enhances performance. The benefits of SPI outweigh the incremental improvements offered by larger backbones, demonstrating that SPI’s impact on accuracy is more substantial than merely increasing the backbone size.
Backbone SPI Self-Supervised Supervised
$AbsRel\downarrow$ $RMSE\downarrow$ $AbsRel\downarrow$ $RMSE\downarrow$
ConvNeXt Large - 0.075 3.722 0.043 1.698
ConvNeXt Large 0.072 3.677 0.037 1.597
ConvNeXt X-Large 0.071 3.670 0.034 1.529
ConvNeXtV2 Huge 0.071 3.662 0.029 1.394
Table 5: Ablation Study Results on KITTI Dataset. The table compares the performance of different backbone networks with and without Strengthened Pose Information (SPI) in both self-supervised and supervised settings.
@@ 6 Conclusion
In summary, SPIdepth achieves significant advancements in self-supervised monocular depth estimation by enhancing the pose network during training, with no changes needed for inference. Despite adding only a minimal number of parameters compared to the depth model, SPIdepth delivers exceptional accuracy.
On the KITTI, Cityscapes, and Make3D datasets, SPIdepth sets new benchmarks in both self-supervised and fine-tuning settings, outperforming models that use multiple frames for inference. Its effectiveness in scenarios with dynamic objects and zero-shot settings demonstrates its robustness and versatility.
These results highlight SPIdepth’s potential for real-world applications, offering precise depth estimation and superior performance across diverse challenges, ready for rapid adaptation and deployment.